FQCoDel(8) Linux FQCoDel(8) NAME CoDel−FairQueuing(FQ)withControlledDelay(CoDel) SYNOPSIS tcqdisc...fqcodel[limitPACKETS][flowsNUMBER][targetTIME][intervalTIME ][quantumBYTES][ecn|noecn] DESCRIPTION FQCodel(FairQueuingControlledDelay)isqueuingdisciplinethatcombinesFairQueuingwith theCoDelAQMscheme.FQCodelusesastochasticmodeltoclassifyincomingpacketsintodifferentflowsandisusedtoprovideafairshareofthebandwidthtoalltheflowsusingthequeue. EachsuchflowismanagedbytheCoDelqueuingdiscipline.Reorderingwithinaflowisavoided sinceCodelinternallyusesaFIFOqueue. PARAMETERS limit hasthesamesemanticsascodelandisthehardlimitontherealqueuesize.Whenthislimitis reached,incomingpacketsaredropped.Defaultis10240packets. flows isthenumberofflowsintowhichtheincomingpacketsareclassified.Duetothestochasticnature ofhashing,multipleflowsmayendupbeinghashedintothesameslot.Newerflowshavepriority overolderones.Thisparametercanbesetonlyatloadtimesincememoryhastobeallocatedfor thehashtable.Defaultvalueis1024. target hasthesamesemanticsascodelandistheacceptableminimumstanding/persistentqueuedelay. Thisminimumdelayisidentifiedbytrackingthelocalminimumqueuedelaythatpacketsexperience.Defaultvalueis5ms. interval hasthesamesemanticsascodelandisusedtoensurethatthemeasuredminimumdelaydoes notbecometoostale.Theminimumdelaymustbeexperiencedinthelastepochoflength.B interval.Itshouldbesetontheorderoftheworst-caseRTTthroughthebottlenecktogiveend-pointssufficienttimetoreact.Defaultvalueis100ms. quantum isthenumberofbytesusedas’deficit’inthefairqueuingalgorithm.Defaultissetto1514bytes whichcorrespondstotheEthernetMTUplusthehardwareheaderlengthof14bytes. ecn|noecn hasthesamesemanticsascodelandcanbeusedtomarkpacketsinsteadofdroppingthem.If ecnhasbeenenabled,noecncanbeusedtoturnitoffandvice-a-versa.Unlikecodel,ecnis turnedonbydefault. EXAMPLES #tcqdiscadddeveth0rootfqcodel #tc-sqdiscshow qdiscfqcodel8002:deveth0rootrefcnt2limit10240pflows1024quantum1514 target5.0msinterval100.0msecn Sent428514bytes2269pkt(dropped0,overlimits0requeues0) backlog0b0prequeues0 iproute2 4June2012 1 FQCoDel(8) Linux FQCoDel(8) maxpacket256dropoverlimit0newflowcount0ecnmark0 newflowslen0oldflowslen0 #tcqdiscadddeveth0rootfqcodellimit2000target3msinterval40msnoecn #tc-sqdiscshow qdiscfqcodel8003:deveth0rootrefcnt2limit2000pflows1024quantum1514target3.0ms interval40.0ms Sent2588985006bytes1783629pkt(dropped0,overlimits0requeues34869) backlog0b0prequeues34869 maxpacket65226dropoverlimit0newflowcount73ecnmark0 newflowslen1oldflowslen3 SEEALSO tc(8),tc-codel(8),tc-red(8) AUTHORS FQCoDelwasimplementedbyEricDumazet.ThismanpagewaswrittenbyVijaySubramanian. PleasereportcorrectionstotheLinuxNetworkingmailinglist<netdev@vger.kernel.org>. iproute2 4June2012 2